solve the D.E. ydouble prime =1+( y prime)^2
1
Expert's answer
2020-03-24T09:41:52-0400

"y''=1+(y')^2"

"y'=dy\/dx=t"

"y''=dt\/dx=t'"

"dt\/dx=1+t^2"

"\int dt/(1+t^2)=\int dx"

"arctan(t)=x+C_1"

"t=tan(x+C_1)"

"dy\/dx=tan(x+C_1)"

"\int dy=\int tan(x+C_1)dx"

substitution for the integral: "z=cos(x+C_1), dz=-sin(x+C_1)dx"

"y=\int -dz/z=-log(z)+C_2=-log(cos(x+C_1))+C_2"

answer: "y=-log(cos(x+C_1))+C_2"